# IWR-Bench Team

### v1

@@ Abstract
The webpage-to-code task requires models to understand visual representations of webpages and generate corresponding code. However, existing benchmarks primarily focus on static screenshot-to-code tasks, thereby overlooking the dynamic interactions fundamental to real-world web applications. To address this limitation, this paper introduces IWR-Bench, a novel benchmark for evaluating the capabilities of Large Vision-Language Models (LVLMs) in interactive webpage reconstruction from video. IWR-Bench comprises 113 meticulously curated tasks from 100 real-world websites, with 1,001 actions and featuring diverse interaction complexities (e.g., web games), visual styles, and domains. Aligning with standard web development practices, each task includes not only user interaction videos but also all crawled static assets (e.g., images, videos). This benchmark evaluates models on two fundamental challenges: comprehensive multi-modal reasoning to infer interaction logic from video and assets, and advanced code generation to translate this logic into functional code. An agent-as-a-judge framework with a comprehensive metric system automatically assesses the functional correctness and visual fidelity of generated webpages. Extensive experiments on 28 LVLMs reveal a significant challenge: the best model achieves an overall score of only 36.35%, as functional correctness (24.39% IFS) lags significantly behind visual fidelity (64.25% VFS). These results highlight critical limitations in current models’ ability to reason about temporal dynamics and synthesize event-driven logic, establishing IWR-Bench as a challenging frontier for vision-language research. The benchmark and evaluation code will be made publicly available. Code is available at https://github.com/L-O-I/IWR-Bench.
@@ 1 Introduction
Figure 2: Overview of the IWR-Bench task and evaluation. The inputs to the model are (a) a user interaction video and (b) composite images of all static assets sniffed from the webpage. The evaluation employs an agent-as-judge framework (Zhuge et al., 2024), where an automated agent assesses the rendered page’s interactivity by executing (c) a ground-truth action sequence and its visual fidelity through screenshot comparison.
@@
Recent advances in Large Vision-Language Models (LVLMs) have unlocked remarkable capabilities in visual understanding and code generation (OpenAI, 2025; Comanici et al., 2025; Bai et al., 2025). State-of-the-art models can now translate a static screenshot of a webpage into corresponding HTML with impressive fidelity (Yun et al., 2024; Gui et al., 2025). This nascent success, however, highlights a fundamental limitation of current evaluation methodologies. Existing benchmarks are either confined to static reconstruction (e.g., Design2Code (Si et al., 2024), WebSight (Laurençon et al., 2024)) or model interactions as single-step, stateless events from image pairs (e.g., Interaction2Code (Xiao et al., 2025)), while also failing to provide the necessary static assets for reconstruction. This simplified setup falls short of capturing the continuous, stateful workflows and complete resource context characteristic of real-world web applications. The disconnect between demonstrated capabilities and the demands of true interactivity motivates our central research question: Can LVLMs reconstruct the dynamic, interactive functionalities of a webpage from observing a user interaction video?
@@
Reconstructing an interactive webpage from video poses two fundamental challenges. The first, comprehensive multi-modal perception and reasoning (Luo et al., 2024; Gupta & Kembhavi, 2023; Song et al., 2025; Deka et al., 2017; Lee et al., 2023), is the process of inferring latent interaction logic from dynamic visual evidence. This requires a model to ground its temporal understanding of observed interactions in a precise visual comprehension of the resultant UI states. A critical facet of this reasoning is robust image matching to associate dynamic elements with their static asset counterparts. The second challenge, advanced code generation (Jimenez et al., 2024; Xiao et al., 2025; Li et al., 2022), is the translation of this inferred logic into functional code that implements the complex, stateful logic of interactive applications (e.g., web-games like 2048 and Minesweeper).
@@
The construction of a comprehensive benchmark for interactive webpage reconstruction confronts three pivotal challenges. The first pertains to ensuring Diverse Interaction Coverage, which necessitates the curation of tasks spanning a broad spectrum of interaction paradigms and visual complexities, while simultaneously adhering to strict standardization for reproducible evaluation. The second challenge centers on the establishment of an Authentic Task Environment. Departing from prior benchmarks characterized by incomplete setups or placeholder assets (Jiang et al., 2025; Gui et al., 2025), this requires the meticulous curation of a complete set of authentic resources from live websites. Such resources must encompass both static assets, such as images and icons, and dynamic content, such as embedded videos, to faithfully represent real-world development contexts. The final challenge lies in the formulation of a Robust Automated Evaluation protocol. Conventional metrics, including pixel-wise similarity, are insufficient for this purpose (Zhang et al., 2018; Caron et al., 2021; Radford et al., 2021), as they cannot appraise functionality. An effective protocol must therefore employ programmatic interaction with the generated webpage to ascertain both the functional integrity of its components and the state-wise visual consistency across dynamic transitions.
@@ 2 Related Work
Webpage Understanding. Webpage understanding evolved from structural analysis based on DOM parsing to a subsequent multimodal perspective that jointly represents a page’s visual and textual content (Furuta et al., 2023; Burns et al., 2023; Liu et al., 2024a). Large Vision-Language Models (LVLMs) have advanced webpage understanding by enabling a unified approach where a single model demonstrates strong performance across diverse downstream tasks, indicative of deep comprehension, such as element grounding (Team, 2025) and screen-based question answering (Wang et al., 2024b; Xu et al., 2024). Among these capabilities, generating code from a visual webpage representation is a key task where existing models have demonstrated strong performance (Beltramelli, 2018; Yun et al., 2024; Gui et al., 2025). With the enhanced capabilities of LVLMs in handling multiple images or videos (Bai et al., 2025; OpenAI, 2025; Guo et al., 2025; Comanici et al., 2025), a logical extension of this capability is the generation of interactive webpages, moving beyond static layouts to better mimic real-world applications.
LVLM Benchmarks. The development of benchmarks for LVLMs has been driven by the rapid expansion of their capabilities, leading to evaluations of increasing complexity (Jimenez et al., 2024; Yang et al., 2024; Mialon et al., 2023; Lu et al., 2023). This progression is evident in the evolution from single-image comprehension to multi-image reasoning and video understanding (Yue et al., 2024; Li et al., 2023; Wang et al., 2024a; Liu et al., 2024b; Hu et al., 2025; Li et al., 2024; Fang et al., 2024; Fu et al., 2025; Ning et al., 2023; Chen et al., 2024; Yang et al., 2024; Lu et al., 2025). Concurrently, in the web domain, benchmarks have targeted either webpage understanding or static code generation from a single screenshot (Beltramelli, 2018; Laurençon et al., 2024; Yun et al., 2024; Si et al., 2024; Gui et al., 2025; Jiang et al., 2025; Awal et al., 2025; Xu et al., 2025), with works like IW-Bench (Guo et al., 2024) creating more robust evaluation metrics for this task. A recent advancement, Interaction2Code (Xiao et al., 2025), extends this by generating code from discrete interaction traces. However, such approaches primarily evaluate single-step, stateless events, rather than the complete, stateful workflows captured in continuous video. Therefore, a critical disconnect exists between model capabilities for dynamic inputs and the benchmarks for interactive web generation.
@@ 3.1 Task Definition and Structure
  • Static Web Assets: All relevant images, icons, and videos necessary for reconstruction. To prevent models from leveraging prior knowledge based on semantic filenames (e.g., logo.png), all asset filenames are anonymized (e.g., renamed to asset_001.png) (Agrawal et al., 2018; Gurari et al., 2018). This forces the model to rely on visual matching and reasoning.
@@ 3.2 Benchmark Construction
Interaction Recording and Asset Collection. For each curated task, interactions on the live website are performed by trained annotators and captured as screen recordings, while a browser extension concurrently records the action type $a_{i}$ and parameters $p_{i}$ for each action (Yun et al., 2024; Zhou et al., 2023). In parallel, all relevant static assets, such as images and icons, are collected via automated crawlers and manual inspection.
@@
Verification. Each annotated task undergoes a two-stage quality assurance process. First, a cross-verification review by a different annotator assesses trajectory correctness, asset completeness, and checkpoint fidelity, with a large model additionally used to verify the accuracy of logical assertions against the ground truth (Zheng et al., 2023; Gou et al., 2025). All identified discrepancies are rectified. Finally, disputed, ambiguous, or overly trivial tasks are filtered out, leaving a final collection of 113 verified tasks.
@@ 3.4 Comparison to Other Benchmarks
Pix2Code (Beltramelli, 2018) Image-to-Code Synthesized 1.7K
DWCG (Yun et al., 2024) Image-to-Code Synthesized 60K
WebSight (Laurençon et al., 2024) Image-to-Code Synthesized 2M
Design2Code (Si et al., 2024) Image-to-Code C4 484
CC-HARD (Gui et al., 2025) Image-to-Code C4 128
ScreenCoder (Jiang et al., 2025) Image-to-Code Live Websites 3K
Interaction2Code (Xiao et al., 2025) Images-to-Code C4 & GitHub 374 ✓(Single-step)
@@
MMBench-Video (Fang et al., 2024) Video QA - 609
MVBench (Li et al., 2024) Video QA - 20K
Video-MME (Fu et al., 2025) Video QA - 900
Video-MMMU (Hu et al., 2025) Video QA - 300
@@ 4.1 Evaluation Protocol
The evaluation of generated code $C$ is conducted using a deterministic executor built upon the browser-use library (browser-use, 2025). This executor programmatically interacts with the rendered webpage by sequentially executing each pre-defined action $a_{i}$ from the ground-truth action trajectory $T$. This design isolates the evaluation to code execution and removes any dependency on high-level task planning, thereby ensuring a stable and reproducible protocol.
@@
For logical assertion verification, an MLLM judge, specifically Gemini-2.5-Pro (Comanici et al., 2025), is employed to analyze screenshots of the page state before and after an action to determine its correctness. The prompt for this judge is detailed in Appendix E. Upon the successful completion of an action, a new screenshot is captured. If the visual evaluation flag $v_{i}$ for this step is true, the new screenshot undergoes a visual fidelity assessment. This assessment is based on a composite score that integrates OCR-based text similarity (Cui et al., 2025), DINO-based structural similarity (Oquab et al., 2023), and a high-level evaluation also conducted by Gemini-2.5-Pro (see Appendix E for the prompt). Actions where $v_{i}$ is false, which typically involve insignificant or stochastic visual changes, are omitted from this visual assessment phase.
@@ 5.2 Main Results
   Doubao-seed-1.6 (bytedance, 2025) 65.95 55.62 60.79 22.55 34.02
   Claude-Sonnet-4 (anthropic, 2025b) 65.75 56.92 61.34 22.29 34.00
   Claude-Opus-4 (anthropic, 2025a) 65.23 55.13 60.18 21.83 33.33
   GPT-5-mini (OpenAI, 2025) 63.36 50.25 56.81 23.18 33.27
   GPT-4.1 (OpenAI, 2025) 63.07 54.63 58.85 20.48 31.99
@@
   Gemini-2.5-Pro (Comanici et al., 2025) 57.46 48.91 53.18 20.51 30.31
   GPT-4o (latest) (Hurst et al., 2024) 63.39 51.71 57.55 17.55 29.55
   Gemini-2.5-Flash (Comanici et al., 2025) 47.53 37.75 42.64 19.88 26.71
   GPT-5-nano (OpenAI, 2025) 53.49 35.70 44.59 18.17 26.10
   Grok-4 (X.ai, 2025) 48.95 30.54 39.74 19.44 25.53
   GPT-4o (0806) (Hurst et al., 2024) 54.03 39.83 46.93 15.87 25.19
   Doubao-seed-1.6-flash (bytedance, 2025) 45.49 32.06 38.78 16.34 23.07
   Gemini-2.5-Flash-Lite (Comanici et al., 2025) 28.95 19.05 24.00 13.29 16.50
@@
   Qwen2.5-VL-72B (Bai et al., 2025) 47.83 28.25 38.04 17.42 23.61
   Qwen2.5-VL-32B (Bai et al., 2025) 39.36 23.30 31.33 16.50 20.95
   Keye-VL-1.5-8B (Yang et al., 2025) 30.81 15.49 23.15 16.06 18.18
   MiniCPM-V-4.5 (Yu et al., 2025) 31.18 15.41 23.29 15.58 17.89
   Qwen2.5-VL-7B (Bai et al., 2025) 28.92 12.20 20.56 13.28 15.47
@@
   Mimo-VL-7B (Team et al., 2025a) 23.28 4.99 14.14 10.57 11.64
   GLM-4.5V (Team et al., 2025c) 16.31 10.52 13.41 10.11 11.10
@@
   VideoLLaMA3-7B (Zhang et al., 2025) 31.29 11.86 21.58 10.29 13.67
   InternVideo-2.5-Chat-8B (Wang et al., 2025) 17.27 3.33 10.30 9.97 10.07
@@ 5.4 Validation of the Evaluation Protocol
The robustness and reliability of the evaluation protocol are validated through a rigorous, two-part analysis that addresses both the metric parameters and the agent-as-a-judge methodology (Zheng et al., 2023; Gou et al., 2025; Maaz et al., 2023). First, the weighting coefficients ($w$ and $\alpha$) for the scoring metrics are determined through a human alignment study, with the detailed procedure and results presented in Appendix C. Second, the agent-as-a-judge framework is validated through a multi-stage process. This process includes a meticulous cross-verification of annotated action trajectories (Section 3.2), automated verification of logical assertions using an MLLM-based judge(Comanici et al., 2025), and manual inspection of the agent’s operational fidelity. For the manual inspection, three PhD students observed the agent’s execution on 100 randomly sampled, model-generated webpages, with the browser’s headless mode disabled to compare on-screen behavior against evaluation logs. Failures in the agent’s evaluation were observed in only three instances. These issues typically stemmed from ambiguous element descriptors (e.g., buttons with identical names), which required a more precise locator (d_i). All identified discrepancies were subsequently rectified.

### v3

@@ Abstract
The webpage-to-code task requires models to understand visual representations of webpages and generate corresponding code. However, existing benchmarks primarily focus on static screenshot-to-code tasks, thereby overlooking the dynamic interactions fundamental to real-world web applications. To address this limitation, this paper introduces IWR-Bench, a novel benchmark for evaluating the capabilities of Large Vision-Language Models (LVLMs) in interactive webpage reconstruction from video. IWR-Bench comprises 113 meticulously curated tasks from 100 real-world websites, with 1,001 actions and featuring diverse interaction complexities (e.g., web games), visual styles, and domains. Aligning with standard web development practices, each task includes not only user interaction videos but also all crawled static assets (e.g., images, videos). This benchmark evaluates models on two fundamental challenges: comprehensive multi-modal reasoning to infer interaction logic from video and assets, and advanced code generation to translate this logic into functional code. An agent-as-a-judge framework with a comprehensive metric system automatically assesses the functional correctness and visual fidelity of generated webpages. Extensive experiments on 28 LVLMs reveal a significant challenge: the best model achieves an overall score of only 36.35%, as functional correctness (24.39% IFS) lags significantly behind visual fidelity (64.25% VFS). These results highlight critical limitations in current models’ ability to reason about temporal dynamics and synthesize event-driven logic, establishing IWR-Bench as a challenging frontier for vision-language research. The benchmark and evaluation code will be made publicly available at https://github.com/SIGMME/IWR-Bench.
@@ 1 Introduction
Figure 2: Overview of the IWR-Bench task and evaluation. The inputs to the model are (a) a user interaction video and (b) composite images of all static assets sniffed from the webpage. The evaluation employs an agent-as-judge framework (zhuge2024agent), where an automated agent assesses the rendered page’s interactivity by executing (c) a ground-truth action sequence and its visual fidelity through screenshot comparison.
@@
Recent advances in Large Vision-Language Models (LVLMs) have unlocked remarkable capabilities in visual understanding and code generation (openai2025gpt; comanici2025gemini; bai2025qwen2). State-of-the-art models can now translate a static screenshot of a webpage into corresponding HTML with impressive fidelity (yun2024web2code; gui2025latcoder). This nascent success, however, highlights a fundamental limitation of current evaluation methodologies. Existing benchmarks are either confined to static reconstruction (e.g., Design2Code (si2024design2code), WebSight (laurenccon2024unlocking)) or model interactions as single-step, stateless events from image pairs (e.g., Interaction2Code (xiao2025interaction2code)), while also failing to provide the necessary static assets for reconstruction. This simplified setup falls short of capturing the continuous, stateful workflows and complete resource context characteristic of real-world web applications. The disconnect between demonstrated capabilities and the demands of true interactivity motivates our central research question: Can LVLMs reconstruct the dynamic, interactive functionalities of a webpage from observing a user interaction video?
@@
Reconstructing an interactive webpage from video poses two fundamental challenges. The first, comprehensive multi-modal perception and reasoning (luo2024layoutllm; gupta2023visual; song2025tianyue; deka2017rico; lee2023pix2struct), is the process of inferring latent interaction logic from dynamic visual evidence. This requires a model to ground its temporal understanding of observed interactions in a precise visual comprehension of the resultant UI states. A critical facet of this reasoning is robust image matching to associate dynamic elements with their static asset counterparts. The second challenge, advanced code generation (jimenez2024swe; xiao2025interaction2code; li2022competition), is the translation of this inferred logic into functional code that implements the complex, stateful logic of interactive applications (e.g., web-games like 2048 and Minesweeper).
@@
The construction of a comprehensive benchmark for interactive webpage reconstruction confronts three pivotal challenges. The first pertains to ensuring Diverse Interaction Coverage, which necessitates the curation of tasks spanning a broad spectrum of interaction paradigms and visual complexities, while simultaneously adhering to strict standardization for reproducible evaluation. The second challenge centers on the establishment of an Authentic Task Environment. Departing from prior benchmarks characterized by incomplete setups or placeholder assets (jiang2025screencoder; gui2025latcoder), this requires the meticulous curation of a complete set of authentic resources from live websites. Such resources must encompass both static assets, such as images and icons, and dynamic content, such as embedded videos, to faithfully represent real-world development contexts. The final challenge lies in the formulation of a Robust Automated Evaluation protocol. Conventional metrics, including pixel-wise similarity, are insufficient for this purpose (zhang2018unreasonable; caron2021emerging; radford2021learning), as they cannot appraise functionality. An effective protocol must therefore employ programmatic interaction with the generated webpage to ascertain both the functional integrity of its components and the state-wise visual consistency across dynamic transitions.
@@ 2 Related Work
Webpage Understanding. Webpage understanding evolved from structural analysis based on DOM parsing to a subsequent multimodal perspective that jointly represents a page’s visual and textual content (furuta2023multimodal; burns2023suite; liu2024visualwebbench). Large Vision-Language Models (LVLMs) have advanced webpage understanding by enabling a unified approach where a single model demonstrates strong performance across diverse downstream tasks, indicative of deep comprehension, such as element grounding (showdown2025) and screen-based question answering (wang2024webquest; xu2024hierarchical). Among these capabilities, generating code from a visual webpage representation is a key task where existing models have demonstrated strong performance (beltramelli2018pix2code; yun2024web2code; gui2025latcoder). With the enhanced capabilities of LVLMs in handling multiple images or videos (bai2025qwen2; openai2025gpt; guo2025seed1; comanici2025gemini), a logical extension of this capability is the generation of interactive webpages, moving beyond static layouts to better mimic real-world applications.
LVLM Benchmarks. The development of benchmarks for LVLMs has been driven by the rapid expansion of their capabilities, leading to evaluations of increasing complexity (jimenez2024swe; yangswe; mialon2023gaia; lu23mathvista). This progression is evident in the evolution from single-image comprehension to multi-image reasoning and video understanding (yue2024mmmu; li2023seedbench2; wang2024muirbench; liu2024mmbench; hu2025video; li2024mvbench; fang2024mmbench; fu2025video; ning2023video; chen2024autoeval; yangswe; lu2025webgenbenchevaluatingllmsgenerating). Concurrently, in the web domain, benchmarks have targeted either webpage understanding or static code generation from a single screenshot (beltramelli2018pix2code; laurenccon2024unlocking; yun2024web2code; si2024design2code; gui2025latcoder; jiang2025screencoder; awal2025webmmu; xu2025web), with works like IW-Bench (guo2024iw) creating more robust evaluation metrics for this task. A recent advancement, Interaction2Code (xiao2025interaction2code), extends this by generating code from discrete interaction traces. However, such approaches primarily evaluate single-step, stateless events, rather than the complete, stateful workflows captured in continuous video. Therefore, a critical disconnect exists between model capabilities for dynamic inputs and the benchmarks for interactive web generation.
@@ 3.1 Task Definition and Structure
  • Static Web Assets: All relevant images, icons, and videos necessary for reconstruction. To prevent models from leveraging prior knowledge based on semantic filenames (e.g., logo.png), all asset filenames are anonymized (e.g., renamed to asset_001.png) (agrawal2018don; gurari2018vizwiz). This forces the model to rely on visual matching and reasoning.
@@ 3.2 Benchmark Construction
Interaction Recording and Asset Collection. For each curated task, interactions on the live website are performed by trained annotators and captured as screen recordings, while a browser extension concurrently records the action type $a_{i}$ and parameters $p_{i}$ for each action (yun2024web2code; zhou2023webarena). In parallel, all relevant static assets, such as images and icons, are collected via automated crawlers and manual inspection.
@@
Verification. Each annotated task undergoes a two-stage quality assurance process. First, a cross-verification review by a different annotator assesses trajectory correctness, asset completeness, and checkpoint fidelity, with a large model additionally used to verify the accuracy of logical assertions against the ground truth (zheng2023judging; gou2025mind2web2evaluatingagentic). All identified discrepancies are rectified. Finally, disputed, ambiguous, or overly trivial tasks are filtered out, leaving a final collection of 113 verified tasks.
@@ 3.4 Comparison to Other Benchmarks
Pix2Code (beltramelli2018pix2code) Image-to-Code Synthesized 1.7K
DWCG (yun2024web2code) Image-to-Code Synthesized 60K
WebSight (laurenccon2024unlocking) Image-to-Code Synthesized 2M
Design2Code (si2024design2code) Image-to-Code C4 484
CC-HARD (gui2025latcoder) Image-to-Code C4 128
ScreenCoder (jiang2025screencoder) Image-to-Code Live Websites 3K
Interaction2Code (xiao2025interaction2code) Images-to-Code C4 & GitHub 374 ✓(Single-step)
@@
MMBench-Video (fang2024mmbench) Video QA - 609
MVBench (li2024mvbench) Video QA - 20K
Video-MME (fu2025video) Video QA - 900
Video-MMMU (hu2025video) Video QA - 300
@@ 4.1 Evaluation Protocol
The evaluation of generated code $C$ is conducted using a deterministic executor built upon the browser-use library (browser-use). This executor programmatically interacts with the rendered webpage by sequentially executing each pre-defined action $a_{i}$ from the ground-truth action trajectory $T$. This design isolates the evaluation to code execution and removes any dependency on high-level task planning, thereby ensuring a stable and reproducible protocol.
@@
For logical assertion verification, an MLLM judge, specifically Gemini-2.5-Pro (comanici2025gemini), is employed to analyze screenshots of the page state before and after an action to determine its correctness. The prompt for this judge is detailed in Appendix E. Upon the successful completion of an action, a new screenshot is captured. If the visual evaluation flag $v_{i}$ for this step is true, the new screenshot undergoes a visual fidelity assessment. This assessment is based on a composite score that integrates OCR-based text similarity (cui2025paddleocr), DINO-based structural similarity (oquab2023dinov2), and a high-level evaluation also conducted by Gemini-2.5-Pro (see Appendix E for the prompt). Actions where $v_{i}$ is false, which typically involve insignificant or stochastic visual changes, are omitted from this visual assessment phase.
@@ 5.2 Main Results
   Doubao-seed-1.6 (bytedancese) 65.95 55.62 60.79 22.55 34.02
   Claude-Sonnet-4 (claudesonnet4) 65.75 56.92 61.34 22.29 34.00
   Claude-Opus-4 (claudeopus) 65.23 55.13 60.18 21.83 33.33
   GPT-5-mini (openai2025gpt) 63.36 50.25 56.81 23.18 33.27
   GPT-4.1 (openaigpt41) 63.07 54.63 58.85 20.48 31.99
@@
   Gemini-2.5-Pro (comanici2025gemini) 57.46 48.91 53.18 20.51 30.31
   GPT-4o (latest) (hurst2024gpt) 63.39 51.71 57.55 17.55 29.55
   Gemini-2.5-Flash (comanici2025gemini) 47.53 37.75 42.64 19.88 26.71
   GPT-5-nano (openai2025gpt) 53.49 35.70 44.59 18.17 26.10
   Grok-4 (grok_4) 48.95 30.54 39.74 19.44 25.53
   GPT-4o (0806) (hurst2024gpt) 54.03 39.83 46.93 15.87 25.19
   Doubao-seed-1.6-flash (bytedancese) 45.49 32.06 38.78 16.34 23.07
   Gemini-2.5-Flash-Lite (comanici2025gemini) 28.95 19.05 24.00 13.29 16.50
@@
   Qwen2.5-VL-72B (bai2025qwen2) 47.83 28.25 38.04 17.42 23.61
   Qwen2.5-VL-32B (bai2025qwen2) 39.36 23.30 31.33 16.50 20.95
   Keye-VL-1.5-8B (yang2025kwaikeyevl15technical) 30.81 15.49 23.15 16.06 18.18
   MiniCPM-V-4.5 (yu2025minicpmv45cookingefficient) 31.18 15.41 23.29 15.58 17.89
   Qwen2.5-VL-7B (bai2025qwen2) 28.92 12.20 20.56 13.28 15.47
@@
   Mimo-VL-7B (coreteam2025mimovltechnicalreport) 23.28 4.99 14.14 10.57 11.64
   GLM-4.5V (vteam2025glm45vglm41vthinkingversatilemultimodal) 16.31 10.52 13.41 10.11 11.10
@@
   VideoLLaMA3-7B (zhang2025videollama) 31.29 11.86 21.58 10.29 13.67
   InternVideo-2.5-Chat-8B (wang2025internvideo2) 17.27 3.33 10.30 9.97 10.07
@@ 5.4 Validation of the Evaluation Protocol
The robustness and reliability of the evaluation protocol are validated through a rigorous, two-part analysis that addresses both the metric parameters and the agent-as-a-judge methodology (zheng2023judging; gou2025mind2web2evaluatingagentic; maaz2023video). First, the weighting coefficients ($w$ and $\alpha$) for the scoring metrics are determined through a human alignment study, with the detailed procedure and results presented in Appendix C. Second, the agent-as-a-judge framework is validated through a multi-stage process. This process includes a meticulous cross-verification of annotated action trajectories (Section 3.2), automated verification of logical assertions using an MLLM-based judge(comanici2025gemini), and manual inspection of the agent’s operational fidelity. For the manual inspection, three PhD students observed the agent’s execution on 100 randomly sampled, model-generated webpages, with the browser’s headless mode disabled to compare on-screen behavior against evaluation logs. Failures in the agent’s evaluation were observed in only three instances. These issues typically stemmed from ambiguous element descriptors (e.g., buttons with identical names), which required a more precise locator (d_i). All identified discrepancies were subsequently rectified.